just 15 not 16, solved 16 already15. 0/5 points | Previcus Answers GiancolIT 26.P062. My Notes Ask Your Teacher An electron mo = 9.11 x 1031 kg is accelerated from rest to speed v by a conservative force. In this process, its potential energy decreases by 6.6 × 10-14·Determine the electrons speed v 4.4e-22X 16. 0/5 points | Previcus Answers Giancolir 26.P:068. My Notes Ask Your Teacher The fictional starship Enterprise obtains its power by combining matter and antimatter, achieving complete conversion of mass into energy. If the mass of the Enterprise is approximately 5.4 x 109 kg, how much mass must be converted into kinetic energy to accelerate it from rest to one-fourth the speed of light? Ignore the decrease in total mass for the calculation. 0.5 x kg
Answer #1

15.)

6.6 * 10^-14 = 1/2 m (vf^2 -vi^2)

6.6 * 10^-14 * 2 / m = vf^2

vf = 3.8 * 10^8 m/s

16.)

Energy gained by the body as it accelerated from rest to 0.25 time of light speed,

E = 0.5*M*v^2

= 0.5*5.4*10^9*(0.25*3*10^8)^2

= 1.518*10^25 J

now let m is the mass that should be converted into energy.

Use Einsteins equation

E = m*c^2

==> m =E/c^2

= (1.518*10^25)/(3*10^8)^2

= 1.68*10^8 kg